A mercury lamp, which emits several discrete wavelengths of light, finds many uses in spectroscopy. Mercury lines are employed as wavelength standards, as UV sources, and historically as excitation sources for Raman scattering. The 435.83-nm light from a mercury lamp is used to irradiate a pure liquid CCl4 sample, and scattered light is observed at 447.57, 442.19, and 440.05 nm (among others.) What CCl4 vibrational wavenumbers correspond to the scattered light observed here?

447.57nm: _____ cm^-1

44219nm: _____ cm^-1

440.05nm: _____ cm^-1
Answer #1

given data are the wave length of scattered light , to convert it to wave numbers in cm^-1 ,first convert unit of wavelengths in

to cm , then calculate 1/wavelength = wave number

a.  447.57nm: =447.57 x10^-7 cm

wave number = 1/447.57 x10^-7 cm = 2.23 x10^-10 cm^-1

b.  44219nm: =44219 x 10^-7 cm

wave number = 1/44219 x10^-7 cm = 2.26 x10^-12 cm^-1

c. 440.05nm: _= 440.05 x10^-7 cm

wave number 1/440.05 x10^-7 cm = 2.27 x10^-10 cm^-1

*********

hope it is clear !